Grade 11Algebra

if a set has 4 elements write how many subsets does it can have?

kavya sharma
9 Years agoGrade 11

2 Answers

eknath
9 Years ago
elements in set A are 4.
total no. of subset = 2^4
                           =2*2*2*2
                           =4*4
                           =16
 total number ob sub set in set A are 16
 
Vikas TU
9 Years ago
No. of subsets of a set = 2n=24=16
No. of proper subsets =2n-1=15.
There’s a differemce in a set and the proper set.